# Low-degree mod 2 cohomology of classifying spaces of $G_{2}$-gauge groups

### v1

@@ Abstract.
In this paper we begin a complementary study of the mod $2$ cohomology of the classifying spaces $B\mathcal{G}_{k}(G_{2})$. Our goal is to understand the structure of $H^{*}(B\mathcal{G}_{k};\mathbb{F}_{2})$ as an unstable module over the mod $2$ Steenrod algebra $\mathcal{A}$ in a low range of degrees. Using the evaluation fibration
@@
from $u_{5}$ to the degree-$6$ generator $x_{6}\in H^{6}(BG_{2};\mathbb{F}_{2})$, for a certain scalar $\varepsilon(k)\in\mathbb{F}_{2}$ depending on the bundle class $k$.
@@ 1.3. Main result: a low-degree description
Theorem 1.1 should be viewed as a structural reduction. It does not claim an explicit formula for $\varepsilon(k)$, nor does it assert that $\varepsilon(k)$ is nontrivial. Kameko’s result shows that the 22–local homotopy type of $\mathcal{G}_{k}(G_{2})$ depends on $k\bmod 4$ through the 22–primary order of the Samelson product $\langle i_{3},1\rangle$ [8, 7]. It is natural to expect that $\varepsilon(k)$ is a function of this 22–primary data; however, determining $\varepsilon(k)$ explicitly requires detailed knowledge of $H^{*}(\Omega^{3}_{0}G_{2})$ and of the connecting map $\partial_{k}$ on mod $2$ cohomology, and we leave this as an open problem.

### v2

@@ Abstract.
In this paper, we begin a complementary study of the mod $2$ cohomology of the classifying spaces $B\mathcal{G}_{k}(G_{2})$. Our goal is to understand the structure of $H^{*}(B\mathcal{G}_{k};\mathbb{F}_{2})$ as an unstable module over the mod $2$ Steenrod algebra in a low range of degrees. Using the evaluation fibration
@@
from $u_{5}$ to the degree-$6$ generator $x_{6}\in H^{6}(BG_{2};\mathbb{F}_{2})$, for a scalar $\varepsilon(k)\in\mathbb{F}_{2}$ encoding the low-degree effect of the bundle class. In addition, 22–locally we prove that $\varepsilon(k)$ is 44–periodic in $k$ (i.e. it depends only on $k\bmod 4$) and that $\varepsilon(k)=0$ for all $k\equiv 0\pmod{4}$.
@@ 1.3. Main result: a low-degree description
Theorem 1.1 should be viewed as a structural reduction. It does not claim a closed formula for $\varepsilon(k)$, nor does it assert that $\varepsilon(k)$ is nontrivial. Kameko’s result shows that the 22–local homotopy type of $\mathcal{G}_{k}(G_{2})$ depends on $k\bmod 4$ through the 22–primary order of the Samelson product $\langle i_{3},1\rangle$ [8, 7]. It is therefore natural to expect that $\varepsilon(k)$ is governed by the same 22–primary data.
In fact, we prove (Lemma 4.5) that, 22–locally, the scalar $\varepsilon(k)$ depends only on the residue class of $k$ modulo $4$, and moreover $\varepsilon(k)=0$ whenever $4\mid k$ (Corollary 4.6). Thus the remaining low-degree problem reduces to determining the two values $\varepsilon(1)$ and $\varepsilon(2)$.
Determining these values explicitly would require a more detailed understanding of $H^{*}(\Omega^{3}_{0}G_{2})$ and of the induced map of the connecting map $\partial_{k}$ on mod $2$ cohomology, and we leave this as a natural problem for future work.
@@ 4.2. Bidegree constraints on differentials
Lemma 4.5 (A 22–local periodicity of $\varepsilon(k)$).
Work 22–locally. If $k\equiv k^{\prime}\pmod{4}$, then
$\varepsilon(k)=\varepsilon(k^{\prime})\in\mathbb{F}_{2}.$
In particular, $\varepsilon(k)$ depends only on the residue class of $k$ modulo $4$.
Proof.
By [8, Lemma 2.1], the connecting map $\partial_{k}\colon G_{2}\to\Omega^{3}_{0}G_{2}$ is the triple adjoint of the Samelson product $\langle k\cdot i_{3},1\rangle$, and in particular $\partial_{k}\simeq k\cdot\partial_{1}$. Kameko proved that the 22–primary component of $\langle i_{3},1\rangle$ has order $4$ [7, Theorem 1.1]. Equivalently, after localization at $2$ the map $\partial_{1}$ has order $4$, so $\partial_{k}\simeq\partial_{k^{\prime}}$ 22–locally whenever $k\equiv k^{\prime}\pmod{4}$.
The Serre spectral sequence (1.1) is functorial with respect to maps of fibrations. Since the kk–dependence of the evaluation fibration is controlled by $\partial_{k}$ (see Section 3), the homotopy $\partial_{k}\simeq\partial_{k^{\prime}}$ induces an identification of the Serre spectral sequences for $k$ and $k^{\prime}$ which is compatible with the standard identifications $E_{2}^{s,t}(k)\cong H^{s}(BG_{2})\otimes H^{t}(\Omega^{3}_{0}G_{2})$. Under this identification, the classes $u_{5}\in E_{2}^{0,5}$ and $x_{6}\in E_{2}^{6,0}\cong H^{6}(BG_{2})$ correspond, and hence the $d_{6}$–values correspond. Since $H^{6}(BG_{2})\cong\mathbb{F}_{2}\{x_{6}\}$ is one-dimensional, the scalar in
$d_{6}(u_{5})=\varepsilon(k)\,x_{6}$
must coincide for $k$ and $k^{\prime}$, proving $\varepsilon(k)=\varepsilon(k^{\prime})$. ∎
Corollary 4.6.
Work 22–locally. If $4\mid k$, then $\varepsilon(k)=0$.
Proof.
If $4\mid k$ then $k\equiv 0\pmod{4}$, so Lemma 4.5 gives $\varepsilon(k)=\varepsilon(0)$. By Note 4.4 we have $\varepsilon(0)=0$, hence $\varepsilon(k)=0$. ∎
Remark 4.7.
Lemma 4.5 and Corollary 4.6 show that, in the low-degree range considered here, the potential kk–dependence through the first fibre class $u_{5}$ reduces to at most three cases:
$k\equiv 0,\ 2,\ 1\text{ or }3\pmod{4}$.
In particular, every class $k\equiv 0\pmod{4}$ falls into the same low-degree behaviour as the trivial bundle, namely $\varepsilon(k)=0$. Thus, to detect genuinely new behaviour via $d_{6}(u_{5})$, it suffices to determine $\varepsilon(1)$ and $\varepsilon(2)$. This reduction is consistent with the 22–local classification of $G_{2}$–gauge groups governed by the 22–primary order of $\langle i_{3},1\rangle$ [8, 7].